40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 3 04:43:03 2011, kiwamu, Update, LSC, sideband-resonance PRMI locked ITMYtable.png
    Reply  Mon Oct 3 13:04:12 2011, Jenne, Update, SUS, AUXEX, AUXEY rebooted 
Message ID: 5600     Entry time: Mon Oct 3 13:04:12 2011     In reply to: 5598
Author: Jenne 
Type: Update 
Category: SUS 
Subject: AUXEX, AUXEY rebooted 

Quote:

  + I found that burtrestore for the ETMX DC coil forces were not functional.

  => currently ETMX's "restore" and "mislalign" buttons on the C1IFO_ALIGN screen are not working.

  => According to the error messages, something seemed wrong on c1auxex, which is a slow machine controlling the DC force.

 [Suresh, Jenne]

Suresh pointed out the oddity that all of the EX, EY slow channels were showing white boxes on the medm screens on all of the workstations except Rosalba.  I don't know why Rosalba seemed to be working, but whatever.  I'm not 100% sure that Rosalba was even working properly....I shutdown ETMX and ETMY's watchdogs before we went to boot computers, but when I came back to the control room the 2 optics were rung up anyway.  Turning back on the watchdogs, the optics immediately began to damp happily.

Since Kiwamu had trouble with the slow channels for EX, we decided to key some crates. 

We keyed the c1auxey, and c1auxex crates, waited a few seconds, and then things looked okay in medm-land.  I looked at the "View Backup" for ETMX, and there were no values for the DC sliders, so since the arms are both flashing right now, I did a "save", and then confirmed that I can misalign and restore the optic.  However, since I didn't fully align/lock the cavity, the saved value for right now shouldn't be fully trusted.  We might have to manually align the cavity using the BS.

ELOG V3.1.3-